# Event Topology Classifiers at the Large Hadron Collider

### v1

@@ II.6 Transverse spherocity ($S_{0}$)
Similar to transverse sphericity, transverse spherocity is another event shape observable extensively used at the LHC energies to separate the soft versus hard QCD-dominated events. Transverse spherocity is defined for a unit vector $\hat{n}(n_{T},0)$ in the transverse plane, which follows a minimization criterion as shown in the following equation [102, 103, 104, 58]
$S_{0}=\frac{\pi^{2}}{4}\min_{\hat{n}}\bigg{(}\frac{\sum_{i=1}^{N_{\rm had}}|\vec{p}_{{\rm T}_{i}}\times\hat{n}|}{\sum_{i=1}^{N_{\rm had}}~{}|\vec{p}_{{\rm T}_{i}}|}\bigg{)}^{2}$ (9)
As previously mentioned, along with the transverse spherocity, most of the discussed observables are infrared and collinear safe quantities. In the next subsections, we discuss the collinear and infrared safety for transverse spherocity as an example.
II.6.1 Infrared Safety
@@ II.6.2 Collinear Safety
Here, $z$ and $(1-z)$ stand for the fractions of the momentum of the hadron in a collinear splitting and $10<z<$ 1 and $\vec{p}_{{\rm T}_{k}}$, $\vec{p}_{{\rm T}_{k1}}$, and $\vec{p}_{{\rm T}_{k2}}$ are parallel to each other. For any direction of $\hat{n}$,
$|\vec{p}_{{\rm T}_{k1}}\times\hat{n}|+|\vec{p}_{{\rm T}_{k2}}\times\hat{n}|=z|\vec{p}_{{\rm T}_{k}}\times\hat{n}|+(1-z)|\vec{p}_{{\rm T}_{k}}\times\hat{n}|=|\vec{p}_{{\rm T}_{k}}\times\hat{n}|.$ (14)
@@
Figure 13: Correlation of $S_{0}$ (left panel) and $S_{0}^{p_{\rm T}=1}$ (right panel) with number of MPI in $pp$ collisions at $\sqrt{s}=13$ TeV using PYTHIA8.
@@
$S_{0}^{p_{\rm T}=1}=\frac{\pi^{2}}{4}\min_{\hat{n}}\Bigg{(}\frac{\sum_{i=1}^{N_{\rm had}}|\hat{p}_{\rm T}\times\hat{n}|}{N_{\rm had}}\Bigg{)}^{2}.$ (16)
@@
Figure 12 shows the distribution of weighted transverse spherocity ($S_{0}$) (left) and unweighted transverse spherocity ($S_{0}^{p_{\rm T}=1}$) (right) for minimum bias $pp$ collisions at $\sqrt{s}=13$ TeV using PYTHIA8. The transverse spherocity distributions are further divided into different percentiles and specified as different transverse spherocity classes. Figure 13 shows the correlation of weighted ($S_{0}$) (left panel) and unweighted ($S_{0}^{p_{\rm T}=1}$) (right panel) transverse spherocity with $N_{\rm mpi}$. The solid line guides the mean values of $N_{\rm mpi}$, i.e., $\langle N_{\rm mpi}\rangle$, for each bin on the $x-$axis. The measurements shown in Fig. 13 are performed in $pp$ collisions at $\sqrt{s}=13$ TeV using PYTHIA8. From Fig. 13, it is conclusive that both $S_{0}$, and $S_{0}^{p_{\rm T}=1}$ retains a fair degree of correlation with $N_{\rm mpi}$, where the higher values of $S_{0}$, and $S_{0}^{p_{\rm T}=1}$ corresponds to a large number of multi-partonic interactions, i.e., soft events. When compared to Fig. 10, the correlation of $S_{\rm T}$ with $N_{\rm mpi}$ appears to be smaller in magnitude as compared to the correlation of transverse spherocity and charged-particle multiplicity with $N_{\rm mpi}$, especially for extreme isotropic events.
@@
Figure 14 shows the correlation of $S_{0}$ and $S_{0}^{p_{\rm T}=1}$ with $N_{\rm mpi}$. A similar correlation of $S_{0}$ and $S_{0}^{p_{\rm T}=1}$ with $N_{\rm mpi}$ is also shown for the highest (0-1)multiplicity class in the mid and forward pseudorapidity regions. Both $S_{0}$ and $S_{0}^{p_{\rm T}=1}$ possess a fair degree of positive correlation with $N_{\rm mpi}$, where events having spherocity value close to one have higher $\langle N_{\rm mpi}\rangle$ value. When no multiplicity selection is applied, $S_{0}$ probes higher values of $\langle N_{\rm mpi}\rangle$ as compared to $S_{0}^{p_{\rm T}=1}$ in the intermediate values of spherocity. The difference of $\langle N_{\rm mpi}\rangle$ for both weighted and unweighted transverse spherocity for the minimum bias case starts to vanish as one goes close to their extreme values, i.e., 0 and 1. In addition, by applying charged-particle event selection, the transverse spherocity achieves a higher value of $\langle N_{\rm mpi}\rangle$. For the top 1% multiplicity class based on both $N_{\rm ch}^{\rm mid}$ and $N_{\rm ch}^{\rm fwd}$ we observe that $S_{0}$ probes a higher value of $N_{\rm mpi}$ than $S_{0}^{p_{\rm T}=1}$ towards lower values of spherocity. As the value of spherocity approaches ‘1’, $\langle N_{\rm mpi}\rangle$ from both $S_{0}^{p_{\rm T}=1}$ and $S_{0}$ approach each other and the difference in $\langle N_{\rm mpi}\rangle$ vanishes for $S_{0}\gtrsim 0.8$. Furthermore, both weighted and unweighted values of transverse spherocity for (0-1)$N_{\rm ch}^{\rm mid}$ class acquired a higher value of $\langle N_{\rm mpi}\rangle$ and a lower slope, relative to (0-1)$N_{\rm ch}^{\rm fwd}$ class. The lower slope of $S_{0}$ and $S_{0}^{p_{\rm T}=1}$ vs $\langle N_{\rm mpi}\rangle$ for the (0-1)class of $N_{\rm ch}^{\rm mid}$, may arise due to estimation of both spherocity and $N_{\rm ch}^{\rm mid}$ in the similar pseudorapidity region.
II.7 Relative transverse activity classifier ($R_{\rm T}$)
@@
In relativistic collisions, the partonic interactions are characterized by pQCD or non-pQCD processes. Initial hard scattering in such collisions is a pQCD process due to the large energy and momentum transfers involved in such collisions. On the other hand, multi-partonic interactions, initial and final state radiations, and beam remnants are non-pQCD in nature. The partons involved in the multi-partonic interactions, initial and final state radiations, and the beam remnant collectively contribute to the underlying event. The particles contributing to UE do not originate from the fragmentation of partons produced in the hardest scattering. Underlying event (UE) activity is one of the efficient tools theorised to affect particle production and can lead to observed QGP-like signatures in $pp$ collisions.
@@
Figure 15: Depiction of different topological regions with respect to the trigger jet particle.
@@
Figure 18: Correlation of $R_{\rm{T}}$ measured in the transverse, trans-min and trans-max regions with number of MPI in $pp$ collisions at $\sqrt{s}=13$ TeV using PYTHIA8.
To classify the events based on the contribution from the UE events, one can define different topological regions based on the contribution from the jet fragmentation and the UE. Figure 15 depicts different topological regions, viz., toward, transverse and away, defined with respect to the highest $p_{\rm T}$ particle, also called the leading particle, and the corresponding momentum is $p_{\rm T}^{\rm leading}$. The toward region, as the name indicates, consists of the particles with an azimuthal angle less than $\pi/3$ relative to the leading particle, i.e., $|\Delta\phi|<\pi/3$ and has the largest contribution from the fragmentation of jets. The away region contains all the charged particles with $|\Delta\phi|\geq 2\pi/3$ while the transverse region is defined to be within $\pi/3\leq|\Delta\phi|<2\pi/3$. Since the transverse region is perpendicular to the leading jet axis, it is expected to have the least contribution from the jet and must be dominated by the underlying event activity. To estimate the value of the relative transverse activity classifier ($R_{\rm T}$), the number of charged hadrons in the transverse regions is determined, denoted as $N_{\rm ch}^{\rm trans}$. Thus, one obtains the value of $R_{\rm T}$ using the following expression.
@@
Similarly, for $R_{\rm T}^{\rm min}$, and $R_{\rm T}^{\rm max}$, the transverse region is again subdivided into $\pi/3<\Delta\phi<2\pi/3$ and $-\pi/3>\Delta\phi>-2\pi/3$. The charged-particle multiplicity in these two regions is determined. For each event, the region with higher multiplicity (trans-max) contributes to $N_{\rm ch}^{\rm T-max}$ while the region with less multiplicity (trans-min) contributes to $N_{\rm ch}^{\rm T-min}$. Thus, one obtains, $R_{\rm T}^{\rm min}$ and $R_{\rm T}^{\rm max}$ as follows [41].
@@
In Equations (17), (18), and (19), the angular brackets represent the event-average value of the observables. Figure 16 shows the distribution of charged-particle yields in the transverse ($N_{\rm ch}^{\rm T}$), trans-min ($N_{\rm ch}^{\rm T-min}$), and trans-max ($N_{\rm ch}^{\rm T-max}$) regions. The charged hadrons are measured in the mid-pseudorapidity region, i.e., $|\eta|<0.8$, with $p_{\rm T}^{\rm lead}>5$ GeV/$c$ in $pp$ collisions at $\sqrt{s}=13$ TeV using PYTHIA8. As expected, in Fig. 16, trans-min region has the lowest multiplicity yield, followed by trans-max. It is important to note that $N_{\rm ch}^{\rm T}=N_{\rm ch}^{\rm T-min}+N_{\rm ch}^{\rm T-max}$, which is also held by the distribution shown in Fig. 16. Figure 17 shows the distribution of $R_{\rm T}$ (upper left), $R_{\rm T}^{\rm min}$ (upper right) and $R_{\rm T}^{\rm max}$ (lower) in $pp$ collisions at $\sqrt{s}=13$ TeV, and corresponding percentile slices.
In Fig. 18, we show the correlation of $R_{\rm T}$ (upper left), $R_{\rm T}^{\rm min}$ (upper right), and $R_{\rm T}^{\rm max}$ (lower) with $N_{\rm mpi}$. Here, the solid line represents the mean values of $N_{\rm mpi}$ for the corresponding bin in the $x-$axis. The measurements shown in Fig. 18 are performed in $pp$ collisions at $\sqrt{s}=13$ TeV using PYTHIA8. All the classifiers, $R_{\rm T}$, $R_{\rm T}^{\rm max}$, and $R_{\rm T}^{\rm min}$ show a fair amount of correlations with $N_{\rm mpi}$. The correlation shown in Fig. 18 is significantly larger as compared to the correlation of transverse spherocity and transverse sphericity with $N_{\rm mpi}$. In other words, using the relative transverse activity classifiers, one can probe to higher values of $N_{\rm mpi}$ compared to $S_{0}$, $S_{0}^{p_{\rm T}=1}$, and $S_{\rm T}$.
II.8 Charged-particle flattenicity ($\rho_{\rm{ch}}$)

### v2

@@ 2.6 Transverse spherocity ($S_{0}$)
Similar to transverse sphericity, transverse spherocity is another event shape observable extensively used at the LHC energies to separate the soft versus hard QCD-dominated events. Transverse spherocity is defined for a unit vector $\hat{n}(n_{T},0)$ in the transverse plane, which follows a minimization criterion as shown in the following equation Cuautle:2014yda , Cuautle:2015kra , Ortiz:2017jho , Khuntia:2018qox
$S_{0}=\frac{\pi^{2}}{4}\min_{\hat{n}}\bigg(\frac{\sum_{i=1}^{N_{\rm had}}|\vec{p}_{{\rm T}_{i}}\times\hat{n}|}{\sum_{i=1}^{N_{\rm had}}~|\vec{p}_{{\rm T}_{i}}|}\bigg)^{2}.$ (11)
Historically, when event shape variables were introduced in $e^{+}e^{-}$ collisions, it was argued that infrared and collinear (IRC) safety of these variables is a key property when one relies on the calculation based on the perturbation series. In other words, event shape variables are used to describe the parton structure in an event, which must be insensitive to the emission of soft and/or collinear radiation. These event shapes include thrust, squared jet-mass, heavy jet mass, jet broadening variables, sphericity, spherocity, aplanarity, etc., which are defined in Sec. 2. Later at the LHC energies, these event shape variables are being widely used for jet-related studies demanding the IRC safety of these quantities, which is a pQCD-based phenomenon. In the next subsections, we discuss the collinear and infrared safety for transverse spherocity as an example. However, it is important to mention that when dealing with the non-perturbative aspects of the collisions using the final state particles such as pions, kaons, etc, the IRC safety of event shape observables is not a key requirement.
2.6.1 Infrared Safety
@@ 2.6.2 Collinear Safety
Here, $z$ and (1z)(1-z) stand for the fractions of the momentum of the hadron in a collinear splitting and $0<z<1$ and $\vec{p}_{{\rm T}_{k}}$, $\vec{p}_{{\rm T}_{k1}}$, and $\vec{p}_{{\rm T}_{k2}}$ are parallel to each other. For any direction of $\hat{n}$,
[…]
@@
Figure 13: Correlation of $S_{0}$ (left panel) and $S_{0}^{p_{\rm T}=1}$ (right panel) with number of MPI in $pp$ collisions at $\sqrt{s}=13$ TeV using PYTHIA8. The black solid line represents $\langle N_{\rm mpi}\rangle$ as a function of $S_{\rm 0}$ (left) and $S_{0}^{p_{\rm T}=1}$ (right).
@@
$S_{0}^{p_{\rm T}=1}=\frac{\pi^{2}}{4}\min_{\hat{n}}\Bigg(\frac{\sum_{i=1}^{N_{\rm had}}|\hat{p}_{{\rm T}_{i}}\times\hat{n}|}{N_{\rm had}}\Bigg)^{2}.$ (18)
@@
Figure 12 shows the distribution of weighted transverse spherocity ($S_{0}$) (left) and unweighted transverse spherocity ($S_{0}^{p_{\rm T}=1}$) (right) for minimum bias $pp$ collisions at $\sqrt{s}=13$ TeV using PYTHIA8. The transverse spherocity distributions are further divided into different percentiles and specified as different transverse spherocity classes. Figure 13 shows the correlation of weighted ($S_{0}$) (left panel) and unweighted ($S_{0}^{p_{\rm T}=1}$) (right panel) transverse spherocity with $N_{\rm mpi}$ obtained in $pp$ collisions at $\sqrt{s}=13$ TeV using PYTHIA8. The solid line guides the mean values of $N_{\rm mpi}$, i.e., $\langle N_{\rm mpi}\rangle$, for each bin on the $x$-axis. From Fig. 13, it is conclusive that both $S_{0}$, and $S_{0}^{p_{\rm T}=1}$ retains a fair degree of correlation with $N_{\rm mpi}$, where the higher values of $S_{0}$, and $S_{0}^{p_{\rm T}=1}$ corresponds to a large number of multi-partonic interactions, i.e., soft events. When compared to Fig. 10, the correlation of $S_{\rm T}$ with $N_{\rm mpi}$ appears to be smaller in magnitude as compared to the correlation of transverse spherocity and charged-particle multiplicity with $N_{\rm mpi}$, especially for extreme isotropic events.
@@
Figure 14 shows the correlation of $S_{0}$ and $S_{0}^{p_{\rm T}=1}$ with $N_{\rm mpi}$. A similar correlation of $S_{0}$ and $S_{0}^{p_{\rm T}=1}$ with $N_{\rm mpi}$ is also shown for the highest (0-1)% multiplicity class in the mid and forward pseudorapidity regions. Both $S_{0}$ and $S_{0}^{p_{\rm T}=1}$ possess a fair degree of positive correlation with $N_{\rm mpi}$, where events having spherocity value close to one have higher $\langle N_{\rm mpi}\rangle$ value. When no multiplicity selection is applied, $S_{0}$ probes higher values of $\langle N_{\rm mpi}\rangle$ as compared to $S_{0}^{p_{\rm T}=1}$ in the intermediate values of spherocity. The difference of $\langle N_{\rm mpi}\rangle$ for both weighted and unweighted transverse spherocity for the minimum bias case starts to vanish as one goes close to their extreme values, i.e., 0 and 1. In addition, by applying charged-particle event selection, the transverse spherocity achieves a higher value of $\langle N_{\rm mpi}\rangle$. For the top 1% multiplicity class based on both $N_{\rm ch}^{\rm mid}$ and $N_{\rm ch}^{\rm fwd}$ we observe that $S_{0}$ probes a higher value of $N_{\rm mpi}$ than $S_{0}^{p_{\rm T}=1}$ towards lower values of spherocity. As the value of spherocity approaches ‘1’, $\langle N_{\rm mpi}\rangle$ from both $S_{0}^{p_{\rm T}=1}$ and $S_{0}$ approach each other and the difference in $\langle N_{\rm mpi}\rangle$ vanishes for $S_{0}\gtrsim 0.8$. Furthermore, both weighted and unweighted values of transverse spherocity for (0-1)% $N_{\rm ch}^{\rm mid}$ class acquired a higher value of $\langle N_{\rm mpi}\rangle$ and a lower slope, relative to (0-1)% $N_{\rm ch}^{\rm fwd}$ class. The lower slope of $S_{0}$ and $S_{0}^{p_{\rm T}=1}$ vs $\langle N_{\rm mpi}\rangle$ for the (0-1)% class of $N_{\rm ch}^{\rm mid}$, may arise due to estimation of both spherocity and $N_{\rm ch}^{\rm mid}$ in the similar pseudorapidity region.
Figure 15: Unfolded distribution of $p_{\rm T}$-weighted transverse spherocity ($S_{0}^{\rm ch}$ or $S_{0}$) in $pp$ collisions at $\sqrt{s}=13$ TeV measured with CMS experiment and compared with predictions from PYTHIA CP1, CP5, A3, A14, tunes, the EPOS-LHC generator, and the HERWIG CH3 tune CMS:2025sws .
Having discussed the correlation of transverse spherocity with $N_{\rm mpi}$, it is observed that the distribution of many event classifiers, including transverse spherocity, from MC event generators fails to reproduce the distribution of event shape observables measured in experiments. A sample is shown in Fig. 15, where the unfolded distribution of transverse spherocity (denoted as $S_{0}^{\rm ch}$) from CMS experiment is compared with MC simulations using PYTHIA CP1 CMS:2019csb , CP5 CMS:2019csb , A3 ATLAS:2016puo , A14 TheATLAScollaboration:2014rfk tunes, the EPOS-LHC generator, and the HERWIG  CMS:2020dqt tune CMS:2025sws . The details about these specific tunes of PYTHIA and HERWIG event generators can be found in appendix 9.1 and 9.3, respectively. The unfolding procedure is used to correct the measured distributions of variables for detector effects, including but not limited to finite resolution and inefficiencies, to retrieve the true underlying physical distribution. This is because the raw data from detectors are usually affected by instrumental limitations. In contrast, unfolding is necessary to make meaningful comparisons with theoretical models or comparisons across experiments. In Fig. 15, it can be observed that the distribution of experimentally measured transverse spherocity is more isotropic in nature than the predictions from different MC event generators. A similar behaviour is observed for other event classifiers, including sphericity, and thrust CMS:2025sws . The comparison of these event shape observables in different multiplicity regions is also made in Ref. CMS:2025sws , which reveals the fact that the mismatch or mis-modeling of event shapes in MC models is likely not coming only from the distribution of the number of charged particles or poorly modeled longitudinal components. These results suggest a need to improve the modeling of $pp$ collisions and are crucial for understanding phenomena such as quark-gluon plasma and topological effects in non-Abelian gauge theories, such as QCD instantons CMS:2025sws .
2.7 Relative transverse activity classifier ($R_{\rm T}$)
@@
In relativistic collisions, the partonic interactions are characterized by pQCD or non-pQCD processes. Initial hard scattering in such collisions is a pQCD process due to the large energy and momentum transfers involved in such collisions. On the other hand, multi-partonic interactions, initial state radiations (ISR), final state radiations (FSR), and beam remnants are non-pQCD in nature. The partons involved in the multi-partonic interactions, initial and final state radiations, and the beam remnant collectively contribute to the underlying event. The particles contributing to UE do not originate from the fragmentation of partons produced in the hardest scattering. UE activity is one of the efficient tools theorised to affect particle production and can lead to observed QGP-like signatures in $pp$ collisions.
@@
Figure 16: Depiction of different topological regions with respect to the leading particle.
@@
Figure 19: Correlation of $R_{\rm{T}}$ measured in the transverse, trans-min and trans-max regions with the number of MPI in $pp$ collisions at $\sqrt{s}=13$ TeV using PYTHIA8. The black solid line represents $\langle N_{\rm mpi}\rangle$ as a function of $R_{\rm T}$ (upper left), $R_{\rm T}^{\rm min}$ (upper right), and $R_{\rm T}^{\rm max}$ (lower middle).
To classify the events based on the contribution from the UE events, one can define different topological regions based on the contribution from the jet fragmentation and the UE. Figure 16 depicts different topological regions, viz., toward, transverse and away, defined with respect to the highest $p_{\rm T}$ particle, also called the leading particle, and the corresponding momentum is $p_{\rm T}^{\rm leading}$. The toward region, as the name indicates, consists of the particles with an azimuthal angle less than $\pi/3$ relative to the leading particle, i.e., $|\Delta\phi|<\pi/3$, and has the largest contribution from the fragmentation of jets. The away region contains all the charged particles with $|\Delta\phi|\geq 2\pi/3$ while the transverse region is defined to be within $\pi/3\leq|\Delta\phi|<2\pi/3$. Since the transverse region is perpendicular to the leading jet axis, it is expected to have the least contribution from the jet and must be dominated by the underlying event activity. To estimate the value of the relative transverse activity classifier ($R_{\rm T}$), the number of charged hadrons in the transverse regions is determined, denoted as $N_{\rm ch}^{\rm trans}$. Thus, one obtains the value of $R_{\rm T}$ using the following expression.
@@
Similarly, for $R_{\rm T}^{\rm min}$, and $R_{\rm T}^{\rm max}$, the transverse region is again subdivided into $\pi/3<\Delta\phi<2\pi/3$ and $\pi/3>\Delta\phi>-2\pi/3$. The charged-particle multiplicity in these two regions is determined. For each event, the region with higher multiplicity (trans-max) contributes to $N_{\rm ch}^{\rm T-max}$ while the region with less multiplicity (trans-min) contributes to $N_{\rm ch}^{\rm T-min}$. Thus, one obtains, $R_{\rm T}^{\rm min}$ and $R_{\rm T}^{\rm max}$ as follows Bencedi:2021tst .
@@
In Equations (19), (20), and (21), the angular brackets represent the event-average value of the observables. Figure 17 shows the distribution of charged-particle yields in the transverse ($N_{\rm ch}^{\rm T}$), trans-min ($N_{\rm ch}^{\rm T-min}$), and trans-max ($N_{\rm ch}^{\rm T-max}$) regions. The charged hadrons are measured in the mid-pseudorapidity region, i.e., $|\eta|<0.8$, with $p_{\rm T}^{\rm lead}>5$ GeV/$c$ in $pp$ collisions at $\sqrt{s}=13$ TeV using PYTHIA8. As expected, in Fig. 17, the trans-min region has the lowest multiplicity yield, followed by trans-max. It is important to note that $N_{\rm ch}^{\rm T}=N_{\rm ch}^{\rm T-min}+N_{\rm ch}^{\rm T-max}$, which is also held by the distribution shown in Fig. 17. Figure 18 shows the distribution of $R_{\rm T}$ (upper left), $R_{\rm T}^{\rm min}$ (upper right) and $R_{\rm T}^{\rm max}$ (lower) in $pp$ collisions at $\sqrt{s}=13$ TeV, and corresponding percentile slices.
In Fig. 19, we show the correlation of $R_{\rm T}$ (upper left), $R_{\rm T}^{\rm min}$ (upper right), and $R_{\rm T}^{\rm max}$ (lower) with $N_{\rm mpi}$ obtained in $pp$ collisions at $\sqrt{s}=13$ TeV using PYTHIA8. Here, the solid line represents the mean values of $N_{\rm mpi}$ for the corresponding bin in the $x$-axis. All the classifiers, $R_{\rm T}$, $R_{\rm T}^{\rm max}$, and $R_{\rm T}^{\rm min}$ show a fair amount of correlations with $N_{\rm mpi}$. The correlation shown in Fig. 19 is significantly larger as compared to the correlation of transverse spherocity and transverse sphericity with $N_{\rm mpi}$. In other words, using the relative transverse activity classifiers, one can probe to higher values of $N_{\rm mpi}$ compared to $S_{0}$, $S_{0}^{p_{\rm T}=1}$, and $S_{\rm T}$.
2.8 Charged-particle flattenicity ($\rho_{\rm{ch}}$)